Groucho has the following test scores: 95%, 91%, 86%, and 92%. What score does he need to earn on his last test to finish with a 91% average?
I've tried to set it up in a proportion, but I keep getting like 300. Can you help?

Found 2 solutions by TimothyLamb, richwmiller:
Answer by TimothyLamb(4379)

(95 + 91 + 86 + 92 + f)/5 = 91
(364 + f)/5 = 91
364 + f = 91*5
364 + f = 455
---
f = 91
---
answer:
he needs 91% on the final to get a class score of 91%
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by richwmiller(17219)

two ways to look at it.
Here's one
let's look at the scores he has so far
95 91 86 92
+4 0 -5 +1
add up the differences =0
He needs 91 because he already has a 91 average
the other tutor gave you the algebra for it